# Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization

### v1

@@ Abstract
Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization
Junjie Lu1thanks:  Equal contribution., Yuliang Liu2,411footnotemark: 1, Chaofeng Qu3, Wei Shen6, Zhouhan Lin2,5, Min Xu1thanks:  Corresponding author 1University of Technology Sydney, 2Shanghai Innovation Institute, 3Southeast University, 4Nanjing University, 5Shanghai Jiao Tong University, 6Independent Researcher {Junjie.Lu,Min.Xu}@uts.edu.au, liuyl03181@gmail.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across various domains, stemming from their ability to reason (Zelikman et al., 2022; Hao et al., 2023; Lewkowycz et al., 2022; Magister et al., 2022; Khalifa et al., 2025). To further enhance their reasoning capabilities, researchers have conducted extensive studies on both the training and inference phases of the models. Training methods typically focus on teaching the model to follow human instruction through supervised training (Yue et al., 2023; Luo et al., 2023), or on preference optimization of reasoning chains using reward models or human feedback, usually known as RLHF (Guo et al., 2025; Zhang et al., 2024a; Le et al., 2022). For inference methods, researchers often employ prompt tuning (Kojima et al., 2022; Imani et al., 2023), or intervene in the model’s generation process by assessing consistency (Wang et al., 2022; Chen et al., 2022; Taubenfeld et al., 2025a), predicting confidence (Weng et al., 2022; Zhu et al., 2025b), and leveraging expected outcomes (Madaan et al., 2023; Cheng et al., 2024).
While chain-of-thought (CoT) (Wei et al., 2023) has become the standard approach for addressing complex reasoning tasks, early training methods primarily focused on outcome optimization, with limited attention to the structure and quality of reasoning paths (Shojaee et al., 2025). One possible is that outcome-based approaches have yielded reasonably satisfactory performance on current benchmarks (Gao et al., 2024b; Hosseini et al., 2024), while another is that assigning fine-grained rewards remains costly. Nevertheless, there have been efforts to optimize finer-grained aspects of the reasoning process, like step-wise methods (Snell et al., 2024; Wang et al., 2024b) and token-wise methods (Zhang et al., 2024b; Lin et al., 2024).
@@
Existing step-wise preference optimization approaches usually locate the model’s first error (Luo et al., 2024; Lu et al., 2024) and construct preference pairs accordingly, enabling focused and targeted optimization. However, these methods often involve substantial labeling overhead, posing a bottleneck in their practical applicability. Furthermore, our analysis of several incorrect answers in mathematical reasoning tasks reveals that models often exhibit confusion, shown by relatively low confidence in generation, well before about 75% of the actual errors occur. Examples are provided in Appendix A.1. This observation inspired us to develop a method capable of identifying issues earlier in a fully automated, human-free manner.
@@
Our experiments show that CGPO yields consistent gains on mathematical reasoning under the same sample counts. The MetaMath-llama model improves by 4.15% on GSM8K (Cobbe et al., 2021) and by 2.54% on MATH (Hendrycks et al., 2021) compared to the baseline. Importantly, our method uses the training data generated by the policy model itself, whereas the baseline uses the training data generated by stronger models or annotated by humans. Models trained with CGPO further demonstrate better generalization on Omni-Math. For the code generation task, we train the DeepSeek-Coder-Instruct-7B by using CGPO and observe a 2.1% improvement on LiveCodeBench (Jain et al., 2024) and 4.0% improvement on LeetCodeDataset (Xia et al., 2025). We also examine the scalability of our method by increasing the training data from 10k to 80k, which results in the data leading to continued improvements.
@@
Figure 2: CGPO overview. The core idea of our method fails in two parts: 1) explore the non-human-like reasoning path by determining the model reasoning step using the initial policy model $\pi_{0}$, and 2) construct chosen/rejected pairs to train $\pi_{0}$ with step-wise DPO training objective.
@@ 2 Related Works
Preference Optimization for LLM: Preference optimization methods have been investigated to improve LLM performance on complex tasks. PPO (Schulman et al., 2017) balances reward maximization with stability through a clipped objective, serving as a foundational method in RLHF but with considerable complexity. Subsequently, critic-free RL methods (Shao et al., 2024; Ahmadian et al., 2024; Hu et al., 2025) reduced complexity by using reward model outputs as immediate rewards for policy gradient updates. And methods like DPO (Rafailov et al., 2024) and SimPO (Meng et al., 2024) directly optimize the model on pairwise preference signals. The reduced computational overhead, combined with sufficiently strong performance, has enabled preference-based methods to achieve widespread success across a variety of domains like instruction following (Wu et al., 2024; Cheng et al., 2025), mathematical reasoning (Pang et al., 2024), and code generation (Zhang et al., 2025; Zhu et al., 2025a).
Confidence-Aware Methods: Model confidence, which is commonly used to quantify model uncertainty, is widely used in LLMs (Fu et al., 2025; Taubenfeld et al., 2025b; Kang et al., 2025; Li et al., 2025; Singhi et al., 2025). While confidence can be interpreted from multiple perspectives. Some approaches measure confidence directly using the predicted probability of the selected token (Farquhar et al., 2024; Liu et al., 2025), which is also the approach adopted in our work. Others rely on the model’s outputs. Xiong et al. (2024) proposed a systematic framework to express model confidence by generation, Yona et al. (2024) measures model confidence by consistency of answers. There are also several methods using the entropy of probability distribution or probe-based methods to measure model confidence in other tasks (Stolfo et al., 2024; Beigi et al., 2024).
Reasoning Trajectory Optimization: LLM reasoning is viewed as a foundational capability for improving performance on complex tasks, and numerous studies have investigated how to improve this ability (OpenAI et al., 2024; Yang et al., 2025): Supervised fine-tuning methods directly align model trajectory to annotated sequence (Yu et al., 2025a); RLHF methods typically force the model to optimize the reasoning trajectory towards a higher score (Yu et al., 2025b); and test time methods intervene in the generation process to optimize reasoning trajectories (Muennighoff et al., 2025). The aforementioned training approaches, especially RLHF, are generally concerned with outcome optimization and do not explicitly enforce correctness along the reasoning trajectories. Lai et al. (2024) directly optimizes an independent reasoning path, the nearest one of our method, while with a significant annotation challenge.
@@ 3 Methods
In this section, we first introduce the basic formulation, then describe the construction of the CGPO
training data, as illustrated in Figure 2. Finally, we introduce the training objective of CGPO
@@ 3.1 Preliminaries
Given an input problem $\mathbf{x}$, an initial policy model $\pi_{0}$ can generate a reasoning sequence under the temperature-based sample strategy. We have $\mathbf{y}=(y_{1},y_{2},\dots,y_{T}),\quad y_{t}\in\mathcal{V}$, where $\mathcal{V}$ denotes the vocabulary. Notably, the sampled token in the generation may not be the token with the top-1 probability. So that at decoding time step $t$, we define the model confidence $c_{t}$ in decoding by the sampled token $y_{t}$, as
$c_{t}\triangleq p(y_{t}|\pi_{0},\mathbf{x},y_{<t})$
@@ 3.2 Preference Pairs Construction
Step Definition. To construct non-human-like step-wise chosen/rejected preference pairs, we first need to clarify the definition of a step in our method. Given $\mathbf{x}$ and the corresponding generated $\mathbf{y}$, we can get the confidence sequence for $\mathbf{y}$, denoted by $\mathbf{c}=(c_{1},c_{2},\dots,c_{T})$. A confidence threshold is employed to split the reasoning steps, with $\tau$ determined based on the distribution of all confidence values $c$ across the dataset. The tokens with confidence below $\tau$ will be the split tokens. Then we get $\mathbf{y}=\mathbf{s}=(s_{1},s_{2},\dots,s_{J})$, where $s_{j}$ denotes a reasoning step that consists of several tokens.
Initial Reasoning Trajectory Determination. To align with the pair-wise preference optimization paradigm, based on the step definition, we choose the sequence before the most uncertain step of $\mathbf{s}$ to serve as the shared initial reasoning trajectory $s_{init}$ of the chosen/rejected pairs.
Chosen/Rejected Pairs Construction. We introduce a reward model $R$ to select the first token of the chosen/rejected pairs given $(\mathbf{x},s_{init})$. The Top-k candidate tokens, ordered by their logits, are denoted as $(v_{1},v_{2},\dots,v_{k})$. The reward for each token is obtained as $R(\mathbf{x},s_{init},v_{k})$. Then, we select the token with the highest score and the one with the lowest score as the starting tokens of the chosen/rejected pairs, denoted by $y^{+}$ and $y^{-}$ respectively.
Subsequently, $\pi_{0}$ will sample the following tokens of $(\mathbf{x},s_{init},y^{+})$ and $(\mathbf{x},s_{init},y^{-})$ before end-of-sequence tag <EOS> or the first token with confidence below $\tau$.
Finally, for each given input, we have two reasoning trajectories: the chosen one $(s_{init},s^{+},\dots)$ and the rejected one $(s_{init},s^{-},\dots)$.
@@ 3.3 Training Objective
Our goal is to optimize $\pi_{\theta}$ based on the triplets $(s_{init},s^{+},s^{-})$. Let $\mathcal{D}$ denote the collection of such triples. With a fixed reference model $\pi_{\text{ref}}$ and temperature $\beta>0$, we optimize $\pi_{\theta}$ using the DPO-like objective
$\mathcal{L}_{\text{CGPO}}(\theta)$ $=-\,\mathbb{E}_{(s_{init},s^{+},s^{-})\sim\mathcal{D}}\Big[\log\sigma\big(\beta(\Delta)\big)\Big],$
$\sigma(a)$ $=\frac{1}{1+e^{-a}}.$
@@
$\Delta$ $=\Delta_{\theta}-\Delta_{\text{ref}},$
$\Delta_{\theta}$ $\triangleq\log\pi_{\theta}(s^{+}\!\mid s_{init})-\log\pi_{\theta}(s^{-}\!\mid s_{init}),$
$\Delta_{\text{ref}}$ $\triangleq\log\pi_{\text{ref}}(s^{+}\!\mid s_{init})-\log\pi_{\text{ref}}(s^{-}\!\mid s_{init}).$
@@
$\log\pi(s^{\pm}\!\mid s_{init})\ =\ \sum_{t}^{seq}\log\pi\!\big(y_{t}\mid x,s_{init}),$
where the sum starts at the branching index $y^{\pm}$. This objective increases the likelihood of the R{R}-selected high-quality trajectory $s^{+}$ while decreasing that of $s^{-}$; the reference term $\Delta_{\text{ref}}$ stabilizes the relative preference signal. We use Algorithm 1 to demonstrate the entire pipeline of CGPO.
@@
1:Dataset $\mathcal{X}$, $\pi_{0}$, $R$, temp, $\tau$, $k$, $\beta$
@@
3:Initialize $\pi_{\theta}\leftarrow\pi_{0}$, preference dataset $\mathcal{D}\leftarrow\emptyset$
@@
5:  Sample $\mathbf{y}$ from $\pi_{0}$; record confidences $c_{t}=\pi_{0}(y_{t}\mid\mathbf{x},\mathbf{y}_{<t})$
6:  Split $\mathbf{y}$ into steps $\mathbf{s}=(s_{1},\dots,s_{J})$ using $\tau$
7:  Find the most uncertain step; set $s_{\text{init}}=\text{concat}(s_{1},\dots,y_{min_{c}})$
8:  Obtain Top-$k$ next-token candidates $(v_{1},\dots,v_{k})$
9:  Score candidates $r_{i}\leftarrow R(\mathbf{x},s_{\text{init}},v_{i})$; choose $y^{+}=\arg\max_{i}r_{i}$ and $y^{-}=\arg\min_{i}r_{i}$
10:  Expand the chosen branch: sample $^{+}, ss^{-}$ from $\pi_{0}$ starting at $^{+}, yy^{-}$, stopping at EOS or the first token with confidence $\tau$
11:  Add $(\mathbf{x},s_{\text{init}},s^{+},s^{-})$ to $\mathcal{D}$
12:end for
13:for $d=1$ to $D$ do
14:  Update $\pi_{\theta}$ by Step-DPO loss, $Loss(\mathbf{x},y^{+}_{d},y^{-}_{d})$
@@
16:return $\pi_{\theta}$

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across various domains, stemming from their ability to reason (Hao et al., 2023; Lewkowycz et al., 2022; Magister et al., 2022; Khalifa et al., 2025). To further enhance their reasoning capabilities, researchers have conducted extensive studies on both the training and inference phases of the models. Training methods typically focus on teaching the model to follow human instruction through supervised training (Yue et al., 2023; Luo et al., 2023), or on preference optimization of reasoning chains using reward models or human feedback, usually known as RLHF (Guo et al., 2025; Zhang et al., 2024a). For inference methods, researchers often employ prompt tuning (Kojima et al., 2022; Imani et al., 2023), or intervene in the model’s generation process by assessing consistency (Wang et al., 2022; Chen et al., 2022; Taubenfeld et al., 2025b), predicting confidence (Weng et al., 2022; Zhu et al., 2025b), and leveraging expected outcomes (Madaan et al., 2023; Cheng et al., 2024).
While chain-of-thought (CoT) (Wei et al., 2023) has become the standard approach for addressing complex reasoning tasks, early training methods primarily focused on outcome optimization, with limited attention to the structure and quality of reasoning paths (Shojaee et al., 2025). One possible reason is that outcome-based approaches have yielded reasonably satisfactory performance on current benchmarks (Gao et al., 2024; Hosseini et al., 2024), while another is that assigning fine-grained rewards remains costly. Nevertheless, there have been efforts to optimize finer-grained aspects of the reasoning process, like step-wise methods (Snell et al., 2024; Wang et al., 2024b) and token-wise methods (Zhang et al., 2024b; Lin et al., 2024).
@@
Existing step-wise preference optimization approaches usually locate the model’s first error (Luo et al., 2024; Lu et al., 2024) and construct preference pairs accordingly, enabling focused and targeted optimization. However, these methods often involve substantial labeling overhead, posing a bottleneck in their practical applicability. Furthermore, our analysis on the incorrect answers of multiple models on mathematical reasoning tasks reveals that models often exhibit confusion, indicated by relatively low confidence in generation, well before about 75% of the actual errors. Examples are provided in Appendix A.1. This observation inspired us to develop a method capable of identifying issues earlier in a fully automated, human-free manner.
@@
Our experiments show that CGPO yields consistent gains on mathematical reasoning under the same number of training samples. Specifically, averaged across all evaluated models, CGPO yields a performance gain of 2.1% on GSM8K (Cobbe et al., 2021) and 1.2% on MATH (Hendrycks et al., 2021) relative to the baseline. Importantly, our method uses the training data generated by the policy model itself, whereas the baseline uses the training data generated by stronger models or annotated by humans. Models trained with CGPO further demonstrate better generalization on CNMO24(Liu et al., 2024). For the code generation task, we train the DeepSeek-Coder-Instruct-7B by CGPO and observe a 0.7% improvement on LiveCodeBench (Jain et al., 2024) and 0.9% improvement on LeetCodeDataset (Xia et al., 2025). We also examine the scalability of our method by increasing the training data from 10k to 80k, which results in the data leading to continued improvements.
@@ 2 Related Works
Preference Optimization for LLM: Preference optimization methods have been investigated to improve LLM performance on complex tasks. PPO (Schulman et al., 2017) balances reward maximization with stability through a clipped objective, serving as a foundational method in RLHF but with considerable complexity, such as involving the critic training and clipping hyperparameter tuning. Subsequently, critic-free RL methods (Shao et al., 2024; Ahmadian et al., 2024; Hu et al., 2025) reduce such complexity by using reward model outputs as immediate rewards for policy gradient updates. The methods like DPO (Rafailov et al., 2024) and SimPO (Meng et al., 2024) directly optimize the model on pairwise preference signals. The reduced computational overhead, combined with sufficiently strong performance, enables preference-based methods to achieve widespread success across a variety of domains like instruction following (Wu et al., 2024; Cheng et al., 2025), mathematical reasoning (Pang et al., 2024), and code generation (Zhang et al., 2025a; Zhu et al., 2025a).
Confidence-Aware Methods: Model confidence, which is commonly used to quantify model uncertainty, is widely used in LLMs (Fu et al., 2025; Taubenfeld et al., 2025a; Kang et al., 2025; Li et al., 2025; Singhi et al., 2025; Xu et al., 2025). Confidence can be interpreted from multiple perspectives. While some approaches measure confidence directly using the prediction probability of selected tokens (Farquhar et al., 2024; Liu et al., 2025), others rely on the model’s outputs. For example, Xiong et al. (2024) proposed a systematic framework to express model confidence by generation, and Yona et al. (2024) measures model confidence by consistency of answers. There are also several methods that use the entropy of probability distribution or probe-based methods to measure model confidence in other tasks (Stolfo et al., 2024; Beigi et al., 2024). In this work, we adopt a token probability-based confidence for its simplicity, generality, and efficiency.
Reasoning Trajectory Optimization: LLM reasoning is viewed as a foundational capability for improving performance on complex tasks, and numerous studies have investigated how to improve this ability (OpenAI et al., 2024; Yang et al., 2025). Supervised fine-tuning methods directly align model trajectory to annotated sequence (Yu et al., 2025a); RLHF methods typically force the model to optimize the reasoning trajectory towards a higher score (Yu et al., 2025b); and test time methods intervene in the generation process to optimize reasoning trajectories (Muennighoff et al., 2025). The aforementioned training approaches, especially RLHF, are generally concerned with outcome optimization and do not explicitly enforce correctness along the reasoning trajectories. In contrast, Lai et al. (2024) directly optimizes an independent reasoning path, the nearest one of our method, while with a significant annotation challenge. Our approach avoids reliance on human-annotated data and instead aligns with the model’s distribution over reasoning trajectories, exhibiting stronger generalization and scalability.
@@ 3 Methods
Figure 2: CGPO overview. The core idea of our method consists of two parts: 1) explore the non-human-like reasoning path by determining the model reasoning step using the initial policy model $\pi_{\theta}$, and 2) construct chosen/rejected pairs to train $\pi_{\theta}$ with step-wise DPO training objective.
In this section, we will introduce the basic formulation, the construction of the training data, and the training objective of CGPO. We present the overall pipeline in Figure 2.
@@ 3.1 Preliminaries
Given a problem $\mathbf{x}$, the initial policy model $\pi_{\theta}$ generates a reasoning sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{T}),\,y_{t}\in\mathcal{V}$, where $\mathcal{V}$ denotes the vocabulary. Notably, the selected token in generation may not be the token with the top-1 probability since we use a temperature-based sampling strategy for exploration. Let $z_{t}(v)$ denote the logit of token $v$ at step $t$. The sampling distribution with temperature $\beta_{\text{smpl}}$ is defined as:
$p_{\beta_{\text{smpl}}}(v\mid\pi_{\theta},\mathbf{x},y_{<t})=\frac{\exp(z_{t}(v)/\beta_{\text{smpl}})}{\sum_{v^{\prime}\in\mathcal{V}}\exp(z_{t}(v^{\prime})/\beta_{\text{smpl}})}.$ (1)
In addition, we define the model confidence $c_{t}$ as:
$c_{t}\triangleq p_{\beta_{\text{smpl}}}(y_{t}\mid\pi_{\theta},\mathbf{x},y_{<t}).$ (2)
@@ 3.2 Preference Pairs Construction
Step Definition. To construct non-human-like step-wise chosen/rejected preference pairs, we first need to define non-human-like steps that differ from those used in previous step-wise optimization methods.
Given $\mathbf{x}$ and the corresponding response $\mathbf{y}$, we can get the confidence sequence for $\mathbf{y}$, denoted by $\mathbf{c}=(c_{1},c_{2},\dots,c_{T})$. A confidence threshold is employed to split the reasoning steps, with $\tau$ determined based on the distribution of all confidence values $c$ across the dataset. We define split points as token positions where confidence falls below the threshold $\tau$, $\mathcal{T}=\{t\mid c_{t}<\tau\}$. Let $\{t_{j}\}_{j=1}^{J}$ be the sorted elements of $\mathcal{T}$, with $t_{0}=1$ and $t_{J+1}=T+1$. Each reasoning step is defined as a contiguous token subsequence:
$\mathbf{s}_{j}:=(y_{t_{j-1}},\dots,y_{t_{j}-1}),\quad j=1,\dots,J+1.$ (3)
Thus, the generated sequence is segmented into steps $\mathbf{s}=(s_{1},\dots,s_{J+1})$.
Initial Reasoning Trajectory Determination. To align with the pair-wise preference optimization paradigm, we select a shared initial reasoning trajectory before an uncertain decision point. Let
$t^{*}=\arg\min_{t\in\{1,\dots,T\}}c_{t}$ (4)
be the index of the least-confident token. We define the shared initial reasoning trajectory as the prefix before $t^{*}$:
$\mathbf{s}_{\text{init}}:=(y_{1},\dots,y_{t^{*}-1}).$ (5)
Chosen/Rejected Pairs Construction. We introduce a reward model $R$ to select the first token of the chosen/rejected pairs given $(\mathbf{x},\mathbf{s}_{init})$. The top-k candidate tokens, ordered by their logits, are denoted as $(v_{1},v_{2},\dots,v_{k})$. The reward for each token is obtained as $R(\mathbf{x},\mathbf{s}_{init},v_{i})$. Then, we select the token with the highest score and the one with the lowest score as the starting tokens of the chosen/rejected pairs, denoted by $y^{+}$ and $y^{-}$ respectively.
Subsequently, $\pi_{\theta}$ samples continuation sequences conditioned on $(\mathbf{x},\mathbf{s}_{\text{init}},y^{+})$ and $(\mathbf{x},\mathbf{s}_{\text{init}},y^{-})$ until the end-of-sequence token or the first token with confidence below $\tau$. Formally, we define
$\mathbf{s}^{+}$ $\sim\pi_{\theta}(\cdot\mid\mathbf{x},\mathbf{s}_{\text{init}},y^{+}),$ (6)
$\mathbf{s}^{-}$ $\sim\pi_{\theta}(\cdot\mid\mathbf{x},\mathbf{s}_{\text{init}},y^{-}),$ (7)
as the chosen and rejected continuation sequences, respectively. Finally, for each given input, we have two reasoning trajectories: the chosen one $(\mathbf{s}_{\text{init}},\mathbf{s}^{+})$ and the rejected one $(\mathbf{s}_{\text{init}},\mathbf{s}^{-})$.
@@ 3.3 Training Objective
Our goal is to optimize $\pi_{\theta}$ based on the triplets $(\mathbf{s}_{init},\mathbf{s}^{+},\mathbf{s}^{-})$. Let $\mathcal{D}$ denote the collection of such triples. With a fixed reference model $\pi_{\text{ref}}$ and temperature $\beta>0$, we optimize $\pi_{\theta}$ using the DPO-like objective
$\mathcal{L}_{\text{CGPO}}(\theta)$ $=-\,\mathbb{E}_{(\mathbf{s}_{init},\mathbf{s}^{+},\mathbf{s}^{-})\sim\mathcal{D}}\Big[\log\sigma\big(\beta(\Delta)\big)\Big],$ (8)
$\sigma(a)$ $=\frac{1}{1+e^{-a}}.$ (9)
@@
$\Delta$ $=\Delta_{\theta}-\Delta_{\text{ref}},$ (10)
$\Delta_{\theta}$ $\triangleq\log\pi_{\theta}(\mathbf{s}^{+}\!\mid\mathbf{s}_{init})-\log\pi_{\theta}(\mathbf{s}^{-}\!\mid\mathbf{s}_{init}),$ (11)
$\Delta_{\text{ref}}$ $\triangleq\log\pi_{\text{ref}}(\mathbf{s}^{+}\!\mid\mathbf{s}_{init})-\log\pi_{\text{ref}}(\mathbf{s}^{-}\!\mid\mathbf{s}_{init}).$ (12)
@@
$\log\pi(s^{\pm}\!\mid\mathbf{s}_{init})\ =\ \sum_{t}^{seq}\log\pi\!\big(y_{t}^{\pm}\mid x,\mathbf{s}_{init}).$ (13)
where the sum starts at the branching token $y^{\pm}$. This objective increases the likelihood of the high-quality trajectory $\mathbf{s}^{+}$ while decreasing that of $\mathbf{s}^{-}$, which are selected by the previously defined reward model $R$; the reference term $\Delta_{\text{ref}}$ stabilizes the relative preference signal. We use Algorithm 1 to demonstrate the entire pipeline of CGPO.
@@
1:Dataset $\mathcal{X}$, $\pi_{\theta}$, $R$, temp, $\tau$, $k$, $\beta$
@@
3:Initialize $\pi_{\theta}$, preference dataset $\mathcal{D}\leftarrow\emptyset$
@@
5:  Sample $\mathbf{y}=(y_{1},\cdots,y_{T})$ from $\pi_{\theta}$
6:  Record confidence $\{c_{t}=\pi_{\theta}(y_{t}\mid\mathbf{x},\mathbf{y}_{<t})\}_{t=1}^{T}$
7:  Split $\mathbf{y}$ into steps $\textbf{s}=(\textbf{s}_{1},\dots,\textbf{s}_{J})$ using $\tau$
8:  Find the most uncertain step $t^{*}:=arg\min_{t}c_{t}$
9:  Set $\textbf{s}_{\text{init}}=(y_{1},\dots,y_{t^{*}})$
10:  Obtain top-$k$ next-token candidates $(v_{1},\dots,v_{k})$
11:  Select $i^{+/-}=arg\max/\min_{i}R(\mathbf{x},s_{\text{init}},v_{i})$
12:  Choose $y^{+}=v_{i^{+}}$ and $y^{-}=v_{i^{-}}$
[…]
14:  Add $(\mathbf{x},\textbf{s}_{\text{init}},\textbf{s}^{+},\textbf{s}^{-})$ to $\mathcal{D}$
@@
16:for each $(\mathbf{s}_{\text{init}},\mathbf{s}^{+},\mathbf{s}^{-})\in\mathcal{D}$ do
17:  Update $\pi_{\theta}$ using the CGPO objective in Eq. (9)
18:end for
19:return $\pi_{\theta}$